OpenStudy (boxman61):

solve: log5(x-2)+log5(x-5)=log5(10)

OpenStudy (aravindg):

do u mean base 5?

OpenStudy (boxman61):

\[\log_{5}(x-2)+\log_{5}(x-5)=\log_{5}10 \]

OpenStudy (anonymous):

x=0 , x=7 ok?

OpenStudy (campbell_st):

use the log/exponential property \[5^{\log _{5}(x)} = x\]

OpenStudy (anonymous):

Boxman reply........

OpenStudy (aravindg):

boxaman tell us if u didnt understand

OpenStudy (anonymous):

yah

OpenStudy (boxman61):

sorry im reading through the book... supposed to solve using the one-to-one property...i know the answer is x=7

OpenStudy (anonymous):

HINT: \[ (x-2)(x-5)=10 \]

OpenStudy (dumbcow):

log(a) + log(b) = log(ab) use this property

OpenStudy (dumbcow):

one to one property if log(a) = log(b) --> then a=b

OpenStudy (boxman61):

thank you all

OpenStudy (anonymous):

Glad to help :)